# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 9.7. Angry Tutte-bagels
The four links injectively correspond to furious edges or vertices or nodes in triangle-strips. Hence it suffices to show that there is no path $rst$ in $O$ with all three nodes $r,s,t$ having triangle-torsos representing links. Assume otherwise. Then two adjacent corners are potter, contradicting 4.7. ∎
@@
We may assume that $O$ traverses the $x(i)$ in the cyclic ordering of $\mathbb{Z}_{4}$. Let $P_{A},P_{B}$ denote the components of $O-x(0)-x(2)$, let $A:=\bigcup_{t\in V(P_{A})}V(G_{t})$ and define $B$ similarly. We use $x(1),x(3)$ to define $P_{C},P_{D}$ and $C,D$ analogously.
@@
Case: $x(0)$ is a furious vertex. So $H_{x(0)}$ is a 4-cycle and $S(0)$ consists of two edges, and we only have to show (ii). The neighbouring torso $H_{t(1)}$ must be 3-connected by Lemma 9.4, so the adhesion-set of $x(0)t(1)$ is included in $A\smallsetminus B$ and avoids $\widehat{S(2)}$. Hence (ii) holds.
Case: $x(0)$ is a furious edge. Then $S(0)$ is the adhesion-set of the edge $x(0)$, and we only have to show (i). Note that $t(1)$ is an end of $x(0)$, so the torso $H_{t(1)}$ is 3-connected. Let $U:=V(H_{t(1)}-S(0))$. Every vertex in $S(0)$ has two neighbours in $G$ that are contained in $U$. To ensure that these neighbours lie in $A\smallsetminus B$, it suffices to show that the adhesion-set $\pi$ of $t(1)t(2)$ is included in $A\smallsetminus B$. Since $t(1)t(2)$ is the first candidate for $x(1)$, with $x(2)$ coming later, $\pi$ can only meet $B$ if $H_{t(2)}$ is a triangle. But then $t(2)$ is the first candidate for $x(1)$, so $x(2)$ comes after $t(2)$. If $H_{t(3)}$ is not a triangle, we are done, and otherwise the tip of the triangle $H_{t(3)}$ avoids $\pi$ by Lemma 9.15. Hence $\pi\subseteq A\smallsetminus B$ as desired.
Case: $x(0)$ is in a triangle-strip. Assume first that $H_{t(1)}$ is 3-connected. Since $H_{t(1)}$ is bad and neighboured by the triangle-torso $H_{x(0)}$, its other neighbouring torso $H_{t(2)}$ must be 3-connected. Hence the edge $t(1)t(2)$ is the first candidate for $x(1)$, and so the edge $t(2)t(3)$ is the first candidate for $x(2)$. Therefore, the entire vertex-set of $H_{t(1)}$ is included in $A\smallsetminus B$, which ensures (i) and (ii).
Assume otherwise that $H_{t(1)}$ is a triangle. Let $\pi$ denote the adhesion-set of the edge $t(1)t(2)$. Note that $\pi$ contains an end of the free edge in $S(0)$ and both vertices in $\pi$ are neighbours of the tip of the triangle $H_{x(0)}$. Hence to show both (i) and (ii), it suffices to show that $\pi$ is included in $A\smallsetminus B$. The node $t(1)$ is the first candidate for $x(1)$, so $t(2)$ is the first candidate for $x(2)$. By Lemma 9.4, $H_{t(2)}$ is either a triangle or 3-connected. If $H_{t(2)}$ is 3-connected, then $\pi$ is included in $A\smallsetminus B$. Otherwise $H_{t(2)}$ is a triangle. Then $H_{x(0)}$, $H_{t(1)}$ and $H_{t(2)}$ are three consecutive triangles. The way we have selected the $x(i)$ then ensures that $x(2)$ comes after $t(2)$. Again, $H_{t(3)}$ is either a triangle or 3-connected. If $H_{t(3)}$ is 3-connected, then $\pi$ is included in $A\smallsetminus B$. Otherwise, if $H_{t(3)}$ is a triangle, then its tip is distinct from the tip of $H_{t(2)}$ by Lemma 9.15, and $t(3)$ is the first candidate for $x(2)$. So $\pi$ is included in $A\smallsetminus B$.
This completes the proof of (i) and (ii). Hence $(A,B)$ and $(C,D)$ are tetra-separations with $S(A,B)=S(0)\cup S(2)$ and $S(C,D)=S(1)\cup S(3)$. The tetra-separations $(A,B)$ and $(C,D)$ cross with all links of size two as the links are precisely the four non-empty sets $S(i)$. ∎
@@
(i) $\Rightarrow$ (ii). Clearly, $G\neq K_{5}$ since $K_{5}$ does not have a tetra-separation. By Lemma 9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$. By 9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. Since $G$ has no totally-nested tetra-separations by assumption, all torsos of $\mathcal{O}$ are bad. We get $\alpha(\mathcal{O})\geqslant 4$ from Lemma 9.51.
(ii) $\Rightarrow$ (i) is Lemma 9.52. ∎
@@ 10. Quasi-5-connected torsos
This section deals with the quasi-5-connected outcome of Theorem 1; see 10.5.
A mixed-separation+ $(A,B)$ of a graph $G$ almost interlaces a star $\sigma$ of mixed-separations+ of $G$ if $(C,D)\leqslant(A,B)$ or $(C,D)\leqslant(B,A)$ for all $(C,D)\in\sigma$. The difference between ‘interlaces’ and ‘almost interlaces’ is ‘$<$’ versus $\leqslant$’ in the definitions.
@@
(Analogue of [Tridecomp, Lemma 2.6.7]) Let $G$ be a 4-connected graph. Let $\sigma$ be a star of mixed-4-separations of $G$ satisfying the matching-condition. Let $(A,B)$ be a separation of the torso $\tau$ of $\sigma$. Then there exists a mixed-separation $(\hat{A},\hat{B})$ of $G$ such that:
@@
The proof is the same as the proof of [Tridecomp, Lemma 2.6.7], with one exception: we replace Lemma 1.2.3 with the matching-condition. ∎
@@
Let $(A,B)$ be a tetra-separation in a 4-connected graph $G$. Let $(C,D)$ be a mixed-4-separation+ in $G$ with $(A,B)\leqslant(C,D)$. Let $(C^{\prime},D^{\prime})$ denote either the left-right-shift or the right-left-shift of $(C,D)$. Then $(A,B)\leqslant(C^{\prime},D^{\prime})$.
@@
Let $G$ be a 4-connected graph. Let $(U,W)$ be a tetra-separation of $G$, and let $(A,B)$ be a mixed-4-separation of $G$ such that $(U,W)\leqslant(A,B)$. Suppose that $|W\cap(A\smallsetminus B)|\geqslant 1$. Let $(A^{\prime},B^{\prime})$ be either the left-right-sift or the right-left-shift of $(A,B)$. Then $(U,W)<(A^{\prime},B^{\prime})$.
@@
We have $(U,W)\leqslant(A^{\prime},B^{\prime})$ by Lemma 10.2. Recall that $A\smallsetminus B\subseteq A^{\prime}\smallsetminus B^{\prime}$ by 5.9. Hence $W\cap(A^{\prime}\smallsetminus B^{\prime})$ is nonempty, so in particular the inclusion $W\supseteq B^{\prime}$ is proper. ∎
@@
By Lemma 10.1, there is a mixed-$k$-separation+ $(\hat{A},\hat{B})$ of $G$ such that
@@
    $|\hat{A}\smallsetminus\hat{B}|\geqslant 2$ and $|\hat{B}\smallsetminus\hat{A}|\geqslant 2$ by (iii); and
@@
    for every $(U,W)\in\sigma$ we have $|W\cap(\hat{A}\smallsetminus\hat{B})|\geqslant 2$ and $|W\cap(\hat{B}\smallsetminus\hat{A})|\geqslant 2$ by (iv).
Since $G$ is 4-connected, $(\hat{A},\hat{B})$ is a mixed-4-separation by Lemma 6.1. Let $(\bar{A},\bar{B})$ denote the left-right-shift of $(\hat{A},\hat{B})$. Then $(\bar{A},\bar{B})$ is a tetra-separation by 5.10. By Lemma 10.3, $(\bar{A},\bar{B})$ interlaces $\sigma$. ∎
@@
By Lemma 8.36, $\tau$ is 4-connected or a $K_{4}$, and we are done in the latter case. By the contrapositive of Lemma 10.4, every 4-separation of $\tau$ has a side of size five. ∎
@@ 11. An Angry Theorem for 4-connectivity
A key step in proving Tutte’s decomposition theorem is to characterise the graphs that are 2-angry: those that are 2-connected and all whose 2-separators are crossed – hence the name. Tutte proved the first angry theorem: the 2-angry graphs are precisely the 3-connected graphs and the cycles [TutteCon]. For a modern proof, see [Tridecomp, Theorem 3.3.4]. An angry theorem for 3-connectivity has been proved using tri-separations [Tridecomp, Theorem 1.1.5]. However, like here, the angry theorem was not used in the proof of the decomposition theorem. Still, angry theorems have recently found an application in group theory: the local version of the 2-angry theorem [Local2sep] was used as a key ingredient in [StallingsNilpotent], and versions for higher connectivity are expected to be needed to continue in the direction of [StallingsNilpotent]. Here we prove an angry theorem for 4-connectivity. A graph $G$ is 4-angry if it is 4-connected and every tetra-separation of $G$ is crossed by a tetra-separation.
@@
Forward implication. Assume that $G$ is 4-angry. If $G$ has no tetra-separation, then (i) by Lemma 2.3. So $G$ has a tetra-separation $(A,B)$, which is crossed by a tetra-separation $(C,D)$ as $G$ is 4-angry. By the (Crossing Lemma). (4.1), all links have the same size $\ell$, and $\ell\in\{0,1,2\}$.
@@
    If $\ell=2$, then (ii) by 9.53 and since $G\neq K_{5}$ by $|G|\geqslant 8$.
Backward implication. We use one of Lemma 2.3, 7.8, 8.41 and 9.53. Here Lemma 2.3 needs $|G|\geqslant 8$. ∎
@@ 12. Vertex-transitive graphs
[…]
@@
[Godsil, Theorem 3.4.2] Every finite vertex-transitive $d$-regular graph is $k$-connected for $k:=\lceil\frac{2}{3}(d+1)\rceil$.
@@
    cycle of $K_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.7);
@@
    cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell$ if the bags along $O$ alternate between being $K_{4}$’s and being 4-cycles, and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.8);
@@
    cycle of triangle-bags of length $\ell$ if all bags are triangles, and every two non-consecutive edges on $O$ have disjoint adhesion-sets (Figure 1.9);
@@
    cycle of $K_{2,2}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{2,2}$’s, all torsos of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure 1.10).
Observation 12.4.
@@
Let $Z$ be the centre of the crossing-diagram, which consists of four vertices by the (Crossing Lemma). (4.1). Note that $G-Z$ has at least four components.
@@
By 4-connectivity of $G$, each of the $\geqslant 4$ vertices of $G-Z$ has neighbourhood equal to $Z$. So, $G$ is 4-regular by vertex-transitivity. In particular, $G-Z$ has exactly four vertices, and $G[Z]$ is independent. So, $G=K_{4,4}$. ∎
@@
Assume for a contradiction that $\mathcal{O}=:(O,\mathcal{G})$ has a good bag $G_{t}$. Then we have the (Block-bagel bag setting). (8.14) with $t=t(0)$, which we assume in the following. By 8.27, $(U,W)$ is a totally-nested tetra-separation of $G$. By Lemma 12.5, the separator $S(U,W)$ consists of four edges, and either $G[U]=K_{4}$ or $G[W]=K_{4}$. It follows with vertex-transitivity that $G$ is 4-regular and that the neighbourhood around each vertex induces $K_{1}\sqcup K_{3}$ in $G$.
@@
Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$, and let $u,v$ be the two vertices in the centre. By Lemma 12.8, every bag of $\mathcal{O}$ is bad. In particular, every bag of $\mathcal{O}$ is a $K_{2}$ or a triangle. Assume for a contradiction that $\mathcal{O}$ has a bag that is a triangle. Since every bag of $\mathcal{O}$ is bad, it follows that all bags of $\mathcal{O}$ are triangles.
@@
By Lemma 12.9, every bag of the block-bagel $\mathcal{O}=(O,\mathcal{G})$ induced by $(A,B)$ and $(C,D)$ is a $K_{2}$. Hence every edge of $O$ is helpful, so $|O|\geqslant 4$ by the (Helpful Lemma). (8.6). Since each vertex of $G$ contained in $\bigcup_{t\in V(O)}G_{t}$ has degree $\leqslant 4$ in $G$, the graph $G$ is 4-regular. Let $u,v$ be the vertices in the centre. By Lemma 8.7 it follows that $|O|=4$ and $uv\notin E(G)$. Then $G$ is an octahedron. ∎

### v2

@@ 9.7. Angry Tutte-bagels
The four links injectively correspond to furious edges or vertices or nodes in triangle-strips. Hence it suffices to show that there is no path rstrst in $O$ with all three nodes $r,s,t$ having triangle-torsos representing links. Assume otherwise. Then two adjacent corners are potter, contradicting ˜4.7. ∎
@@
We may assume that $O$ traverses the $x(i)$ in the cyclic ordering of $\mathbb{Z}_{4}$. Let $P_{A},P_{B}$ denote the components of Ox(0)x(2)O-x(0)-x(2), let $A:=\bigcup_{t\in V(P_{A})}V(G_{t})$ and define $B$ similarly. We use $x(1),x(3)$ to define $P_{C},P_{D}$ and $C,D$ analogously.
@@
Case: $x(0)$ is a furious vertex. So $H_{x(0)}$ is a 4-cycle and $S(0)$ consists of two edges, and we only have to show ˜(ii). The neighbouring torso $H_{t(1)}$ must be 3-connected by Lemma˜9.4, so the adhesion-set of $x(0)t(1)$ is included in $A\smallsetminus B$ and avoids $\widehat{S(2)}$. Hence ˜(ii) holds.
Case: $x(0)$ is a furious edge. Then $S(0)$ is the adhesion-set of the edge $x(0)$, and we only have to show ˜(i). Note that $t(1)$ is an end of $x(0)$, so the torso $H_{t(1)}$ is 3-connected. Let $U:=V(H_{t(1)}-S(0))$. Every vertex in $S(0)$ has two neighbours in $G$ that are contained in $U$. To ensure that these neighbours lie in $A\smallsetminus B$, it suffices to show that the adhesion-set $\pi$ of $t(1)t(2)$ is included in $A\smallsetminus B$. Since $t(1)t(2)$ is the first candidate for $x(1)$, with $x(2)$ coming later, $\pi$ can only meet $B$ if $H_{t(2)}$ is a triangle. But then $t(2)$ is the first candidate for $x(1)$, so $x(2)$ comes after $t(2)$. If $H_{t(3)}$ is not a triangle, we are done, and otherwise the tip of the triangle $H_{t(3)}$ avoids $\pi$ by Lemma˜9.15. Hence $\pi\subseteq A\smallsetminus B$ as desired.
Case: $x(0)$ is in a triangle-strip. Assume first that $H_{t(1)}$ is 3-connected. Since $H_{t(1)}$ is bad and neighboured by the triangle-torso $H_{x(0)}$, its other neighbouring torso $H_{t(2)}$ must be 3-connected. Hence the edge $t(1)t(2)$ is the first candidate for $x(1)$, and so the edge $t(2)t(3)$ is the first candidate for $x(2)$. Therefore, the entire vertex-set of $H_{t(1)}$ is included in $A\smallsetminus B$, which ensures ˜(i) and ˜(ii).
Assume otherwise that $H_{t(1)}$ is a triangle. Let $\pi$ denote the adhesion-set of the edge $t(1)t(2)$. Note that $\pi$ contains an end of the free edge in $S(0)$ and both vertices in $\pi$ are neighbours of the tip of the triangle $H_{x(0)}$. Hence to show both ˜(i) and ˜(ii), it suffices to show that $\pi$ is included in $A\smallsetminus B$. The node $t(1)$ is the first candidate for $x(1)$, so $t(2)$ is the first candidate for $x(2)$. By Lemma˜9.4, $H_{t(2)}$ is either a triangle or 3-connected. If $H_{t(2)}$ is 3-connected, then $\pi$ is included in $A\smallsetminus B$. Otherwise $H_{t(2)}$ is a triangle. Then $H_{x(0)}$, $H_{t(1)}$ and $H_{t(2)}$ are three consecutive triangles. The way we have selected the $x(i)$ then ensures that $x(2)$ comes after $t(2)$. Again, $H_{t(3)}$ is either a triangle or 3-connected. If $H_{t(3)}$ is 3-connected, then $\pi$ is included in $A\smallsetminus B$. Otherwise, if $H_{t(3)}$ is a triangle, then its tip is distinct from the tip of $H_{t(2)}$ by Lemma˜9.15, and $t(3)$ is the first candidate for $x(2)$. So $\pi$ is included in $A\smallsetminus B$.
This completes the proof of ˜(i) and ˜(ii). Hence $(A,B)$ and $(C,D)$ are tetra-separations with $S(A,B)=S(0)\cup S(2)$ and $S(C,D)=S(1)\cup S(3)$. The tetra-separations $(A,B)$ and $(C,D)$ cross with all links of size two as the links are precisely the four non-empty sets $S(i)$. ∎
@@
˜(i) $\Rightarrow$ ˜(ii). Clearly, $G\neq K_{5}$ since $K_{5}$ does not have a tetra-separation. By Lemma˜9.10, $(A,B)$ and $(C,D)$ induce a Tutte-bagel $\mathcal{O}$. By ˜9.41, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. Since $G$ has no totally-nested tetra-separations by assumption, all torsos of $\mathcal{O}$ are bad. We get $\alpha(\mathcal{O})\geqslant 4$ from Lemma˜9.51.
˜(ii) $\Rightarrow$ ˜(i) is Lemma˜9.52. ∎
@@ 10. Quasi-5-connected torsos
This section deals with the quasi-5-connected outcome of Theorem˜1; see ˜10.5.
A mixed-separation+ $(A,B)$ of a graph $G$ almost interlaces a star $\sigma$ of mixed-separations+ of $G$ if $(C,D)\leqslant(A,B)$ or $(C,D)\leqslant(B,A)$ for all $(C,D)\in\sigma$. The difference between ‘interlaces’ and ‘almost interlaces’ is ‘<<’ versus $\leqslant$’ in the definitions.
@@
[…]
@@
The proof is the same as the proof of [CK23, Lemma 2.6.7], with one exception: we replace Lemma 1.2.3 with the matching-condition. ∎
@@
Let $(A,B)$ be a tetra-separation in a 4-connected graph $G$. Let $(C,D)$ be a mixed-4-separation+ in $G$ with $(A,B)\leqslant(C,D)$. Let $(C^{\prime},D^{\prime})$ denote either the left-right-reduction or the right-left-reduction of $(C,D)$. Then $(A,B)\leqslant(C^{\prime},D^{\prime})$.
@@
Let $G$ be a 4-connected graph. Let $(U,W)$ be a tetra-separation of $G$, and let $(A,B)$ be a mixed-4-separation of $G$ such that $(U,W)\leqslant(A,B)$. Suppose that $|W\cap(A\smallsetminus B)|\geqslant 1$. Let $(A^{\prime},B^{\prime})$ be either the left-right-sift or the right-left-reduction of $(A,B)$. Then $(U,W)<(A^{\prime},B^{\prime})$.
@@
We have $(U,W)\leqslant(A^{\prime},B^{\prime})$ by Lemma˜10.2. Recall that $A\smallsetminus B\subseteq A^{\prime}\smallsetminus B^{\prime}$ by ˜5.9. Hence $W\cap(A^{\prime}\smallsetminus B^{\prime})$ is nonempty, so in particular the inclusion $W\supseteq B^{\prime}$ is proper. ∎
@@
[…]
@@
    $|\hat{A}\smallsetminus\hat{B}|\geqslant 2$ and $|\hat{B}\smallsetminus\hat{A}|\geqslant 2$ by ˜(iii); and
@@
    for every $(U,W)\in\sigma$ we have $|W\cap(\hat{A}\smallsetminus\hat{B})|\geqslant 2$ and $|W\cap(\hat{B}\smallsetminus\hat{A})|\geqslant 2$ by ˜(iv).
[…]
@@
By Lemma˜8.36, $\tau$ is 4-connected or a $K_{4}$, and we are done in the latter case. By the contrapositive of Lemma˜10.4, every 4-separation of $\tau$ has a side of size five. ∎
@@ 11. An Angry Theorem for 4-connectivity
A key step in proving Tutte’s decomposition theorem is to characterise the graphs that are 2-angry: those that are 2-connected and all whose 2-separators are crossed – hence the name. Tutte proved the first angry theorem: the 2-angry graphs are precisely the 3-connected graphs and the cycles [T66]. For a modern proof, see [CK23, Theorem 3.3.4]. An angry theorem for 3-connectivity has been proved using tri-separations [CK23, Theorem 1.1.5]. However, like here, the angry theorem was not used in the proof of the decomposition theorem. Still, angry theorems have recently found an application in group theory: the local version of the 2-angry theorem [C22] was used as a key ingredient in [CKKT24], and versions for higher connectivity are expected to be needed to continue in the direction of [CKKT24]. Here we prove an angry theorem for 4-connectivity. A graph $G$ is 4-angry if it is 4-connected and every tetra-separation of $G$ is crossed by a tetra-separation.
@@
Forward implication. Assume that $G$ is 4-angry. If $G$ has no tetra-separation, then ˜(i) by Lemma˜2.3. So $G$ has a tetra-separation $(A,B)$, which is crossed by a tetra-separation $(C,D)$ as $G$ is 4-angry. By the (Crossing Lemma). (4.1), all links have the same size $\ell$, and $\ell\in\{0,1,2\}$.
@@
    If $\ell=2$, then ˜(ii) by ˜9.53 and since $G\neq K_{5}$ by $|G|\geqslant 8$.
Backward implication. We use one of Lemma˜2.3, ˜7.8, ˜8.41 and ˜9.53. Here Lemma˜2.3 needs $|G|\geqslant 8$. ∎
[…]
Lemma 12.1.
Let $G$ be a quasi-4-connected graph. Let $(A^{\Delta},B^{\Delta})$ be a $k$-separation of $G^{\Delta}$. Let $U$ be the set of all degree-three vertices of $G$. We obtain $A^{s}$ from $A^{\Delta}$ by adding every vertex $u\in U$ with $\Delta(u)\subseteq G^{\Delta}[A^{\Delta}]$. We obtain $B^{s}$ from $B^{\Delta}$ by adding every vertex $u\in U$ with $\Delta(u)\subseteq G^{\Delta}[B^{\Delta}]$ but $\Delta(u)\not\subseteq G^{\Delta}[A^{\Delta}]$. Then $(A^{s},B^{s})$ is a $k$-separation of $G^{s}$ with $A^{s}\cap V(G^{\Delta})=A^{\Delta}$ and $B^{s}\cap V(G^{\Delta})=B^{\Delta}$ and $A^{s}\cap B^{s}=A^{\Delta}\cap B^{\Delta}$, see [CK23, Lemma 2.6.4]. Let $A:=A^{s}\cap V(G)$ and $B:=B^{s}\cap V(G)$. Then $(A,B)$ is a mixed-separation of $G$ of order $\leqslant k$.∎
Lemma 12.2.
Let $G$ be a quasi-4-connected graph on $\geqslant 7$ vertices. Then $G^{\Delta}$ is 4-connected.
Proof.
Assume for a contradiction that $G^{\Delta}$ is not 4-connected. Let $(A^{\Delta},B^{\Delta})$ be a separation of $G^{\Delta}$ of least possible order $k$, so $k\leqslant 3$. From $(A^{\Delta},B^{\Delta})$ we obtain a mixed-separation $(A,B)$ of $G$ of order $\leqslant k$ as in Lemma˜12.1.
We claim that $A\smallsetminus B$ and $B\smallsetminus A$ have size $\geqslant 2$. It suffices to show that $|A\smallsetminus B|\geqslant 2$ (the proof for $|B\smallsetminus A|\geqslant 2$ is a subproof of that for $|A\smallsetminus B|\geqslant 2$). Since $(A^{\Delta},B^{\Delta})$ is proper, there is a vertex $x\in A^{\Delta}\smallsetminus B^{\Delta}$. We distinguish two cases.
Assume first that $x$ is a subdividing vertex of an edge $e=uv$ of $G$. Consider the triangles $\Delta(u),\Delta(v)\subseteq G^{\Delta}$ that were added around $u$ and $v$, respectively. Both contain the vertex $x$, so both are included in $G^{\Delta}[A^{\Delta}]$ but not in $G^{\Delta}[B^{\Delta}]$. Hence both $u$ and $v$ lie in $A^{s}\smallsetminus B^{s}$. Since both $u$ and $v$ are in $G$, they also lie in $A\smallsetminus B$. Hence $|A\smallsetminus B|\geqslant 2$.
Assume otherwise that $x$ is a vertex in $G^{\Delta}$ that is also a vertex in $G$. Then $x$ is not contained in $U$, so $x$ has degree $\geqslant 4$ in $G$. Now $x\in A^{s}\smallsetminus B^{s}$, and $x\in A\smallsetminus B$. Since $(A,B)$ has order $\leqslant 3$, at least one of the four neighbours of $x$ in $G$ also lies in $A\smallsetminus B$. Hence $|A\smallsetminus B|\geqslant 2$.
Therefore, both $A\smallsetminus B$ and $B\smallsetminus A$ have size $\geqslant 2$. Hence $G$ is not quasi-4-connected by Lemma˜5.11, a contradiction. ∎
Example 12.3.
The number 7 in Lemma˜12.2 is optimal. To see this, let $G$ be a cycle of $C_{4}$-bags of length 3. Then $G^{\Delta}$ is a generalised 4-wheel, which is 3-connected but not quasi-4-connected.
@@ 13. Vertex-transitive graphs
In this section, we prove Figure˜1.8. Recall that an $H$-expansion of a $|V(H)|$-regular graph $G$ is a graph obtained from $G$ by replacing the vertices $v$ of $G$ with pairwise disjoint copies $H_{v}$ of $H$ and by replacing every edge $e=uv$ of $G$ with an $H_{u}$$H_{v}$ edge $\hat{e}$ so that $\hat{e}_{1}$ and $\hat{e}_{2}$ share no ends for every two $e_{1}\neq e_{2}\in E(G)$. Thus, the edges $\hat{e}$ form a perfect matching of the $H$-expansion. Note that if $H$ is complete then the $H$-expansion of $G$ is unique up to isomorphism. The clique-expansion of a graph $G$ is the graph $G^{\prime}$ that is obtained from $G$ by replacing every vertex $v\in V(G)$ by a clique $C_{v}$ of size the degree of $v$, and every edge $e=uv$ by a $C_{u}$$C_{v}$ edge $\hat{e}$ such that these edges $\hat{e}$ form a perfect matching in $G^{\prime}$.
@@
[GR01, Theorem 3.4.2] Every finite vertex-transitive $d$-regular graph is $k$-connected for $k:=\lceil\frac{2}{3}(d+1)\rceil$.
@@
    cycle of $K_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.9);
@@
    cycle alternating between $K_{4}$-bags and $C_{4}$-bags of length $\ell$ if the bags along $O$ alternate between being $K_{4}$’s and being 4-cycles, and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.10);
@@
    cycle of triangle-bags of length $\ell$ if all bags are triangles, and every two non-consecutive edges on $O$ have disjoint adhesion-sets (Figure˜1.12);
@@
    cycle alternating between $K_{2,2}$-bags and $C_{4}$-torsos of length $\ell$ if the torsos along $O$ alternate between being $K_{4}$’s and $C_{4}$’s, and every adhesion-set of $\mathcal{O}$ is independent and all adhesion-sets of $\mathcal{O}$ are disjoint (so the $K_{4}$-torsos have $K_{2,2}$-bags) (Figure˜1.11);
  • $\cdot$
    cycle of $C_{4}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $C_{4}$’s and all adhesion-sets of $\mathcal{O}$ induce disjoint $K_{2}$’s (Figure˜1.13);
  • $\cdot$
    cycle of $K_{2,2}$-bags of length $\ell$ if all bags of $\mathcal{O}$ are $K_{2,2}$’s, all torsos of $\mathcal{O}$ are $K_{4}$’s and all adhesion-sets of $\mathcal{O}$ are disjoint (Figure˜1.14).
Observation 13.4.
@@
Let $Z$ be the centre of the crossing-diagram, which consists of four vertices by the (Crossing Lemma). (4.1). Note that GZG-Z has at least four components.
@@
By 4-connectivity of $G$, each of the $\geqslant 4$ vertices of GZG-Z has neighbourhood equal to $Z$. So, $G$ is 4-regular by vertex-transitivity. In particular, GZG-Z has exactly four vertices, and $G[Z]$ is independent. So, $G=K_{4,4}$. ∎
@@
Assume for a contradiction that $\mathcal{O}=:(O,\mathcal{G})$ has a good bag $G_{t}$. Then we have the (Block-bagel bag setting). (8.14) with $t=t(0)$, which we assume in the following. By ˜8.27, $(U,W)$ is a totally-nested tetra-separation of $G$. By Lemma˜13.5, the separator $S(U,W)$ consists of four edges, and either $G[U]=K_{4}$ or $G[W]=K_{4}$. It follows with vertex-transitivity that $G$ is 4-regular and that the neighbourhood around each vertex induces $K_{1}\sqcup K_{3}$ in $G$.
@@
Write $\mathcal{O}=:(O,(G_{t})_{t\in V(O)})$, and let $u,v$ be the two vertices in the centre. By Lemma˜13.8, every bag of $\mathcal{O}$ is bad. In particular, every bag of $\mathcal{O}$ is a $K_{2}$ or a triangle. Assume for a contradiction that $\mathcal{O}$ has a bag that is a triangle. Since every bag of $\mathcal{O}$ is bad, it follows that all bags of $\mathcal{O}$ are triangles.
@@
By Lemma˜13.9, every bag of the block-bagel $\mathcal{O}=(O,\mathcal{G})$ induced by $(A,B)$ and $(C,D)$ is a $K_{2}$. Hence every edge of $O$ is helpful, so $|O|\geqslant 4$ by the (Helpful Lemma). (8.6). Since each vertex of $G$ contained in $\bigcup_{t\in V(O)}G_{t}$ has degree $\leqslant 4$ in $G$, the graph $G$ is 4-regular. Let $u,v$ be the vertices in the centre. By Lemma˜8.7 it follows that $|O|=4$ and $uv\notin E(G)$. Then $G$ is an octahedron. ∎